5 moles of an ideal gas expand isohermally and reversibly from pressure of 10atm to 2atm at 300K. what is largest mass [approx] which can be lifted through hieght 1 m in this expansion?

Grade:11

1 Answers

Vikas TU
7 years ago
Work done isothermally in reversible process is given as:
W = – nRT * log(P1/P2)
put all the values,
W = – 5R*300*log(10/2) = > 20 KJ (approx.)
W = F x displacement = > mgd where d = 1 m.
W = 20,000 = 10m
m = 2000/1000 = 2 kg.